OpenStudy (andijo76):

does anyone know anything about exponents the question is use the laws of exponents and simplify (2^9/10)^2/7

OpenStudy (andijo76):

i will try to draw it out if this is hard to understand

OpenStudy (andijo76):

okay so i wont draw it this thing wont let me

OpenStudy (anonymous):

http://hwhelpguy.weebly.com/ I have a video about Laws of Exponents, see if that helps

OpenStudy (anonymous):

Say you have 3^2 it would be 3*3=9 So for your problem it would be, 2*2*2*2*2*2*2*2*2=512 Your other exponent, you have to solve in ( ) first. Your equation now is (512/10)^2/7 Divide, then add exponent than divide by 7 Hope that helps..

OpenStudy (anonymous):

51.2*51.2=2621(originally 2621.44 but you round it) 2621/7=374(originally 374.428571) From simplifying it, the answer should be 374